A board ten feet long is cut into three pieces. One piece is one foot longer than the shortest piece and two feet shorter than the longest piece. How long is each piece?
Answer by ptaylor(2198)

Let x=the one piece
Then x-1=shortest piece
And 10-(x+(x-1))=10-x-x+1=10-2x+1=longest piece
Sooooo
(10-2x+1)-2=x
10-2x-1=x collect like terms
3x=9
x=3 feet ---the piece
3-1=2 feet=shortest piece
10-6+1=5 feet=longest piece
Hope this helps----ptaylor